The following information concerns production in the Baking Department for March. All direct materials are placed in process at the beginning of production. ACCOUNT Work in Process—Baking Department ACCOUNT NO. Date Item Debit Credit Balance Debit Credit March 1 Bal., 6,000 units, 4/5 completed 20,400 31 Direct materials, 108,000 units 183,600 204,000 31 Direct labor 47,710 251,710 31 Factory overhead 26,840 278,550 31 Goods finished, 109,500 units 269,640 8,910 31 Bal. ? units, 2/5 completed 8,910 a. Determine the number of units in work in process inventory at March 31. units b. Determine the equivalent units of production for direct materials and conversion costs in March. If an amount is zero, enter in "0". Baking Department Equivalent Units of Production for Direct Materials and Conversion Costs For March Whole Units Direct Materials Equivalent Units Conversion Equivalent Units Inventory in process, March 1 Started and completed in March Transferred to finished goods in March Inventory in process, March 31 Total
Answer #1

a. Determine the number of units in work in process inventory at March 31. units

Ending work in process = 6000+108000-109500 = 4500 Units

b. Determine the equivalent units of production for direct materials and conversion costs in March

Whole Unit Direct Materials Equivalent Units Conversion Equivalent Units
Inventory in process, March 1 6000 0 6000/5 = 1200
Started and completed in March 103500 103500 103500
Transferred to finished goods in March 109500 103500 104700
Inventory in process, March 31 4500 4500 4500*2/5 = 1800
Total 114000 108000 106500